Calculate the energy transferred in bulb X in 3Seconds if the current in the circuit is 2 Amperes


1
Expert's answer
2021-05-23T18:57:00-0400

Gives

Time(t)=3sec

Current (i)=2Amp

Resistance=x ohm

Power


"P=i^2X"

Energy transfer in 💡(bulb)

"P=\frac{E}{t}"

"E=P\times t"

"E=i^2Xt"

Put values

"E=2^2\times3X"

"E=12XJ"